OpenStudy (anonymous):

Subtract then simplify by collecting like radical terms if possible 4 √27-3√3

OpenStudy (saifoo.khan):

\[\frac{4}{\sqrt{27} - 3 \sqrt 3}\]

OpenStudy (saifoo.khan):

This is the question right?

OpenStudy (anonymous):

It wants it to be solved then simplify it into radical terms if possible

OpenStudy (saifoo.khan):

This is not possible.

OpenStudy (anonymous):

can u please help me solve mi problum

OpenStudy (saifoo.khan):

This is not possible.

OpenStudy (anonymous):

yeah what is it?

OpenStudy (anonymous):

to saifoo.khan

OpenStudy (anonymous):

I think he means 4 (√9 x√3) - 3 √3 = 4 (3√3) - 3√3 = 12 √3 - 3√3 = 9√3

OpenStudy (anonymous):

what is the like radical term for?

OpenStudy (anonymous):

This is what is confusing me

OpenStudy (anonymous):

It's that you can't add or subtract unline radicals like √27 and √3, so you have to make them alike. You break up √27 into √9 x √3. √9 simplifies to 3 and now you have like radicals in the form of √3 for both and they can now be subtracted.

OpenStudy (anonymous):

Thank you so much that is a big help.

OpenStudy (anonymous):

4 √27 - 3√3 = 4 √ 3^2 * 3 - 3 √3 = 4 * 3 √3 - 3√3 = 12√3 -3 √3 = 9√3